# Unconditional correctness of recent quantum algorithms for factoring and computing discrete logarithms

### v1

@@ 2.1. Notation
We refer the reader to [16] for a detailed textbook on quantum computing, and [4, Chapter 2] for a brief overview.
@@ 2.2. Restricting to a convenient subgroup of $\widehat{(\mathbb{Z}/N\mathbb{Z})^{\times}}$
The goal of this preliminary section is to define a subgroup of the character group of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ that does not have too many elements of small order. This will be crucial for Section 2.4, to reduce the influence of potential counterexamples to the Generalised Riemann Hypothesis. For example, we need to avoid having many characters $\chi_{1},\ldots,\chi_{k}$ modulo $N$ such that $\chi_{i}^{2}=\psi$ for all $i$, where $\psi$ is an exceptional character (in the sense that $L(s,\psi)$ has a zero very close to the line $\operatorname{Re}s=1$). The definition of the suitable subgroup depends on the precise structure of $(\mathbb{Z}/N\mathbb{Z})^{\times}$.
Notation 2.1.
If $G$ is a finite abelian group (written multiplicatively) and $M\geqslant 1$, we write $G^{M}:=\{x^{M}:x\in G\}$ for the subgroup of $M$th powers in $G$ [^7] In particular, $G^{M}$ is not the $M$-fold direct product of $G$.
@@
Let $N>2$, $d:=\lceil\!\sqrt{\log N}\rceil$, $G:=(\mathbb{Z}/N\mathbb{Z})^{\times}$ and let ${M_{*}}\geqslant 1$ be the integer defined in the statement of Lemma 2.4.
@@
which is $\leqslant h^{d/10}$ by Lemma 2.4. ∎
@@ 2.3. Lattice point counting via characters
To prove Theorem 2.18, we will need to study the following quantities.
@@
Observe that $\psi\big{(}b_{i}^{Me_{i}}\big{)}=\iota(\psi)(b_{i}^{e_{i}})$, where $\iota$ is the isomorphism $\widehat{G^{M}}\xrightarrow{\sim}\widehat{G}^{M}$ defined in Lemma 2.2. We may thus rewrite this equality as
@@
$\sum_{\begin{subarray}{c}(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\\ ~\max_{i}|e_{i}|\leqslant H\end{subarray}}\mathbf{1}_{\prod_{i}\mathbf{b}_{i}^{Me_{i}}\equiv 1\,(\mathrm{mod}\,N)}=\frac{1}{\big{\lvert}\widehat{G}^{M}\big{\rvert}}\sum_{\chi\in\widehat{G}^{M}}\sum_{\begin{subarray}{c}(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\\ ~\max_{i}|e_{i}|\leqslant H\end{subarray}}\prod_{i=1}^{d}\chi^{e_{i}}(b_{i})=\frac{1}{\big{\lvert}\widehat{G}^{M}\big{\rvert}}\sum_{\chi\in\widehat{G}^{M}}F_{\chi,H}(b_{1},\ldots,b_{d})$
@@
It is fairly straightforward to estimate $F_{\chi,H}(b_{1},\ldots,b_{d})$ when $\chi$ is a character of small order in $\widehat{G}$. The contribution of these small-order characters gives the expected main term for 3.
@@
where the first equality follows from Lemma 2.2 and the second from the canonical isomorphism $H_{1}^{\perp}\cong\widehat{G_{1}/H_{1}}$ for any subgroup $H_{1}$ of a finite abelian group $G_{1}$.
Clearly, if $\chi\in\mathopen{}\mathclose{{}\left\langle b_{1},\ldots,b_{d}}\right\rangle^{\perp}\cap\widehat{G}^{M}$, we have $F_{\chi,H}(\underline{b})=(2H+1)^{d}$. Hence, by 5, those characters contribute
@@
Note that $C_{I,(z_{i})}$ is either the empty set, or a coset of $\mathopen{}\mathclose{{}\left\langle b_{1},\ldots,b_{d}}\right\rangle^{\perp}\cap\widehat{G}^{M}$ in $\widehat{G}^{M}$. Moreover, if $\chi$ has order $\leqslant e^{10d}$, this set $C_{I,(z_{i})}$ can only be non-empty if all $z_{i}$ are roots of unity of order $\leqslant e^{10d}$, and there are $\leqslant e^{20d}$ such roots of unity. We conclude that
@@
Thus, we can bound the right-hand side of 6 by
@@
where we used that $H\geqslant e^{30d}$ in the last step. By 5, this means that
@@ 2.4. Bounding the contribution of large-order characters
Proposition 2.9 is a standard consequence of the explicit formula for $L(s,\chi)$ and is proved in Appendix A. The logarithmic factors can be somewhat improved, but such refinements are irrelevant here.
The Generalised Riemann Hypothesis is the claim that, for every Dirichlet character $\chi$, $L(s,\chi)$ has no zero $\rho$ with $\frac{1}{2}<\operatorname{Re}\rho<1$. Assuming GRH, Proposition 2.9 thus implies almost square-root cancellation for character sums over primes. Proposition 2.10 will serve as an unconditional substitute for GRH.
@@
This is [11, Theorem 1 (1.7)]. ∎
Our goal is to control the total contribution of all large-order characters in 4. We will do so in Proposition 2.14 (restricting to a suitable sugbroup of $\widehat{G}$). We first prove the following $L^{2}$ bound for a single large-order character.
@@
By Proposition 2.10, the number of Dirichlet characters modulo $N$ whose associated $L$-function has a zero in the region
[…]
is
@@
If a non-principal character $\psi$ modulo $N$ has no zero in the region 7, then by Proposition 2.9 we have
(9) $\mathbb{E}\mathopen{}\mathclose{{}\left[\psi(\mathbf{b}_{1})}\right]\ll X^{-1/(10d)}(\log(NX))^{3}\ll d^{-100}d^{6}\ll d^{-15}.$
@@
We can use 9 to bound the term $\mathbb{E}\mathopen{}\mathclose{{}\left[\chi^{h}(\mathbf{b}_{1})}\right]$, unless $\chi^{h}$ is principal or $L(s,\chi^{h})$ has a zero in the region 7. By 8, the number of values of $h$ with $|h|\leqslant 2H$ for which one of these two situations occurs is
@@
By 9, we deduce that, for all but $O(e^{-d}H)$ values of $|h|\leqslant 2H$, the bound $\mathbb{E}\mathopen{}\mathclose{{}\left[\chi^{h}(\mathbf{b}_{1})}\right]\ll d^{-15}$ holds. Hence
$\mathbb{E}\mathopen{}\mathclose{{}\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}\right]\leqslant(2H+1)^{d}\Big{(}O(e^{-d}H)+O(d^{-15}H)\Big{)}^{d}\ll d^{-10d}H^{2d}$
@@
Lemma 2.11 applies to all characters $\chi$ of order $\geqslant e^{10d}$, but gives a relatively weak upper bound. In Lemma 2.13, we will prove that a stronger bound can be obtained if a small set of exceptions is allowed. We begin by describing the exceptional characters in the following lemma.
@@
Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma 2.11. Let $H\geqslant e^{10d}$.
@@
Then there are integers $-2H\leqslant h_{1},h_{2}\leqslant 2H$ with $0<h_{2}-h_{1}\leqslant e^{3t}$ such that both $L(s,\chi^{h_{1}})$ and $L(s,\chi^{h_{2}})$ have a zero in the region
[…]
@@
As in the proof of Lemma 2.11, we have
@@
Let $I_{1}$ be the set of all $-2H\leqslant h\leqslant 2H$ such that $\chi^{h}$ is principal. Since $\chi$ has order $\geqslant e^{10d}$, we have
@@
Let $I_{2}$ be the set of all $-2H\leqslant h\leqslant 2H$ such that $L(s,\chi^{h})$ has a zero in the region 10. By contradiction, suppose that the conclusion of Lemma 2.12 does not hold. Then any sub-interval of $[-2H,2H]$ of length $e^{3t}$ contains at most one element of $I_{2}$. This implies that
@@
Moreover, for any integer $h\in[-2H,2H]\setminus(I_{1}\cup I_{2})$, the character $\chi^{h}$ is non-principal and has no zero in the region 10, which by Proposition 2.9 implies that
@@
We wish to use a zero-density estimate again to show that there are few characters satisfying the conclusion of Lemma 2.12. For this step to work, we need to restrict to the subgroup $\widehat{G}^{{M_{*}}}$ of the full character group $\widehat{G}$, where ${M_{*}}$ is the integer defined in Lemma 2.4.
@@
Let $t_{0}\geqslant 1$ be the constant from Lemma 2.12. Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma 2.11. Let $H\geqslant e^{10d}$ and let ${M_{*}}\geqslant 1$ be the integer defined in Lemma 2.4.
@@
(12) |$\mathopen{}\mathclose{{}\left|\big{\{}\chi\in\widehat{G}^{{M_{*}}}\ :\ \mathrm{ord}(\chi)\geqslant e^{10d},\ \mathbb{E}\mathopen{}\mathclose{{}\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}\right]>e^{-2td}H^{2d}\big{\}}}\right|\ll e^{td/2}.$

### v2

@@ 1.6. Notation
We refer the reader to [CNbook] for a detailed textbook on quantum computing, and [postquantum, Chapter 2] for a brief overview.
Acknowledgements
2. An illustrative special case
In this section, we prove a weak version of our main theorem in a simplified setting. This special case can be used to obtain an unconditional factoring algorithm for integers $N$ of a special form: RSA moduli that are products of two “safe primes”. [^7] Note that this restrictive assumption is generally not satisfied by RSA moduli used in practice.
Definition 2.1 (RSA moduli with safe primes).
Let $\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$ be the set of integers $N$ of the form $N=P\cdot Q$, where $P,Q\geqslant N^{1/4}$ are distinct primes such that (P1)/2(P-1)/2 and (Q1)/2(Q-1)/2 are also prime.
For $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$, the multiplicative group $(\mathbb{Z}/N\mathbb{Z})^{\times}$ has a very convenient structure, allowing us to give a relatively short proof of the following proposition.
Proposition 2.2.
Let $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$ be a sufficiently large integer. Let $d:=\lceil\!\sqrt{\log N}\rceil$ and $X:=d^{d}$.
Let $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$ be i.i.d. random variables, each uniformly distributed in the set of primes less than $X$. Let $\mathbf{b}_{0}$ be a random variable uniformly distributed in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, independent from the $\mathbf{b}_{i}$’s.
Let $x\in(\mathbb{Z}/N\mathbb{Z})^{\times}$. Then, with probability $\gg 1$, there are integers $0\leqslant e_{0},\ldots,e_{d}\leqslant e^{O(d)}$ such that
$\prod_{i=0}^{d}\mathbf{b}_{i}^{e_{i}}\equiv x\pmod{N}$.
Applying Proposition˜2.2 (with $x$ being a non-trivial square root of $1$ modulo $N$) shows that Regev’s factoring algorithm, with random parameters $\mathbf{b}_{0},\ldots,\mathbf{b}_{d}$ as in that proposition [^8] Again, the presence of one potentially large parameter $\mathbf{b}_{0}$ circumvents the subgroup obstructions discussed in the introduction, while not significantly affecting the efficiency of the quantum algorithm., admits an unconditional proof of correctness for integers $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$.
It should be noted that Theorem˜3.18 not only removes the restriction $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$, it also provides a stronger geometric conclusion required for applications to the discrete logarithm problem.
Notation 2.3.
For the rest of this section, we fix $N$, $d$, $X$, $\mathbf{b}_{0},\ldots,\mathbf{b}_{d}$ and $x$ as in the statement of Proposition˜2.2. In particular, $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$ is assumed to be sufficiently large. Moreover, there are primes $P$, $Q$, $P^{\prime}$ and $Q^{\prime}$ such that $N=P\cdot Q$, $P=2P^{\prime}+1$ and $Q=2Q^{\prime}+1$. Hence,
(3) $G:=(\mathbb{Z}/N\mathbb{Z})^{\times}\cong(\mathbb{Z}/2\mathbb{Z})^{2}\times\mathbb{Z}/P^{\prime}\mathbb{Z}\times\mathbb{Z}/Q^{\prime}\mathbb{Z}\cong\widehat{G}$.
To prove Proposition˜2.2, we require the following lemma from analytic number theory, that ultimately relies on a zero-density estimate for Dirichlet $L$-functions.
Lemma 2.4.
For every $j\geqslant 0$, the number of characters $\psi\in\widehat{G}$ such that
$\Big\lvert\!\operatorname*{{}\mathbb{E}}_{p\leqslant X}\psi(p)\Big\rvert\geqslant e^{-j}$
is at most $e^{O(d)}N^{O(j/\log X)}$. Here, $\operatorname*{{}\mathbb{E}}_{p\leqslant X}$ denotes the average over primes $p\leqslant X$.
We omit the proof of Lemma˜2.4, as it is somewhat tangential to the aim of this explanatory section. [^9] This lemma is a consequence of Propositions 3.9 and 3.10; see Section 3.3 and Appendix A for further details.
Proof of Proposition˜2.2.
Let $H=\lfloor e^{Cd}\rfloor$, where $C$ is a large absolute constant to be chosen later.
We will show that, with probability $\gg 1$, there exist integers $0\leqslant h_{1},\ldots,h_{d}<H$ such that
$x\equiv\mathbf{b}_{0}\mathbf{b}_{1}^{h_{1}}\cdots\mathbf{b}_{d}^{h_{d}}\pmod{N}$,
which clearly implies the desired conclusion. By orthogonality of characters modulo $N$, we have
(4) $\sum_{0\leqslant h_{1},\ldots,h_{d}<H}\mathbf{1}_{\mathbf{b}_{0}\mathbf{b}_{1}^{h_{1}}\cdots\mathbf{b}_{d}^{h_{d}}\equiv x\,(\mathrm{mod}\,N)}=\frac{1}{\varphi(N)}\sum_{\chi\in\widehat{G}}\chi(\mathbf{b}_{0}x^{-1})\prod_{i=1}^{d}\sum_{0\leqslant h<H}\chi^{h}(\mathbf{b}_{i}).$
Let $\chi_{0}$ be the trivial character, and $\chi_{1},\chi_{2},\chi_{3}$ be the three characters of order $2$ in $\widehat{G}$. Each of these four characters $\chi_{j}$ satisfies $\chi_{j}(\mathbf{b}_{i})\in\{\pm 1\}$ for all $i$, so the term $\sum_{0\leqslant h<H}\chi_{j}^{h}(\mathbf{b}_{i})$ is either equal to $H$ or lies in $\{0,1\}$. Hence, the contribution of these four characters to ˜4 is
(5) $\frac{H^{d}}{\varphi(N)}+\frac{1}{\varphi(N)}\sum_{j=1}^{3}\chi_{j}(\mathbf{b}_{0}x^{-1})\,C_{j}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d}).$
for some quantities $C_{j}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})\geqslant 0$. Since $\mathbf{b}_{0}$ is uniformly distributed in $G$, by ˜3 we have
$\mathbb{P}\mathopen{}\mathclose{{\left(\chi_{1}(\mathbf{b}_{0}x^{-1})=\chi_{2}(\mathbf{b}_{0}x^{-1})=\chi_{3}(\mathbf{b}_{0}x^{-1})=1}}\right)=\frac{1}{4},$
as this is the probability that $\mathbf{b}_{0}x^{-1}$ lies in the subgroup of squares modulo $N$, which has index $4$ in $G$. Therefore, with probability $\geqslant 1/4$, the expression ˜5 is bounded below by $H^{d}/\varphi(N)$.
[…]
(6) $\operatorname*{{}\mathbb{E}}\!\big[\mathopen{}\mathclose{{\left|\mathbf{z}}}\right|^{2}\big]=o\big(H^{2d}\big),$
where $\mathbf{z}$ is the random variable defined by
$\mathbf{z}:=\sum_{\begin{subarray}{c}\chi\in\widehat{G}\\ \mathrm{ord}(\chi)>2\end{subarray}}\Bigg\lvert\prod_{i=1}^{d}\sum_{0\leqslant h<H}\chi^{h}(\mathbf{b}_{i})\Bigg\rvert$.
By the triangle inequality for the $L^{2}$-norm and the definition of the random variables $\mathbf{b}_{i}$, we have
$\operatorname*{{}\mathbb{E}}\!\big[\mathopen{}\mathclose{{\left|\mathbf{z}}}\right|^{2}\big]^{1/2}\leqslant\sum_{\begin{subarray}{c}\chi\in\widehat{G}\\ \mathrm{ord}(\chi)>2\end{subarray}}\Bigg(\operatorname*{{}\mathbb{E}}\Bigg[\bigg\lvert\prod_{i=1}^{d}\sum_{0\leqslant h<H}\chi^{h}(\mathbf{b}_{i})\bigg\rvert^{2}\Bigg]\Bigg)^{1/2}=\sum_{\begin{subarray}{c}\chi\in\widehat{G}\\ \mathrm{ord}(\chi)>2\end{subarray}}\Bigg(\operatorname*{{}\mathbb{E}}_{p\leqslant X}\bigg\lvert\!\sum_{0\leqslant h<H}\chi^{h}(p)\bigg\rvert^{2}\Bigg)^{d/2}.$
Expanding the square and exchanging the order of summation, we obtain
$\operatorname*{{}\mathbb{E}}_{p\leqslant X}\bigg\lvert\!\sum_{0\leqslant h<H}\chi^{h}(p)\bigg\rvert^{2}\leqslant\sum_{0\leqslant h_{1},h_{2}<H}\,\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h_{1}-h_{2}}(p)}}\right|\leqslant H\sum_{|h|<H}\,\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h}(p)}}\right|.$
Therefore,
(7) $\operatorname*{{}\mathbb{E}}\!\big[\mathopen{}\mathclose{{\left|\mathbf{z}}}\right|^{2}\big]^{1/2}\leqslant H^{d/2}\sum_{\begin{subarray}{c}\chi\in\widehat{G}\\ \mathrm{ord}(\chi)>2\end{subarray}}\Bigg(\sum_{|h|<H}\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h}(p)}}\right|\Bigg)^{d/2}.$
To estimate the right-hand side of ˜7, we partition the set of characters as follows. For $j\geqslant 0$, let
[…]
(8) $\mathopen{}\mathclose{{\left|E_{j}}}\right|\ll He^{O(d)}N^{O(j/\log X)}\ll He^{O(d)}e^{O(jd/\log d)}$
for all $j\geqslant 0$. [^10] It is here that we crucially relied on the particular structure of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ for $N\in\mathcal{M}_{\scriptscriptstyle{\mathrm{RSA}}}$. These arguments can be extended somewhat, but additional ideas are required to handle the general case (the most difficult case being when $N$ has many prime factors).
Let $j\geqslant 0$ and let $\chi\in E_{j}$. By Lemma˜2.4, there are at most $e^{O(d)}$ integers $0\leqslant|h|<H$ such that $\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h}(p)}}\right|>d^{-2}$. By definition of $E_{j}$, all other integers $0<|h|<H$ satisfy
$\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h}(p)}}\right|\leqslant\min(d^{-2},e^{-j}).$
Hence,
$\Bigg(\sum_{|h|<H}\mathopen{}\mathclose{{\left|\operatorname*{{}\mathbb{E}}_{p\leqslant X}\chi^{h}(p)}}\right|\Bigg)^{d/2}\leqslant\big(e^{O(d)}+2H\min(d^{-2},e^{-j})\big)^{d/2}\leqslant e^{O(d^{2})}+e^{O(d)}H^{d/2}\min(d^{-d},e^{-jd/2}).$
Summing this over all $\chi\in E_{j}$ and $j\geqslant 0$, we can bound ˜7 by
$\operatorname*{{}\mathbb{E}}\!\big[\mathopen{}\mathclose{{\left|\mathbf{z}}}\right|^{2}\big]^{1/2}\leqslant e^{O(d^{2})}\varphi(N)H^{d/2}+e^{O(d)}H^{d}\sum_{j\geqslant 0}\mathopen{}\mathclose{{\left|E_{j}}}\right|\min(d^{-d},e^{-jd/2}).$
Note that $\varphi(N)\leqslant N\leqslant e^{d^{2}}$. Bounding $\mathopen{}\mathclose{{\left|E_{j}}}\right|$ using ˜8 and summing the resulting geometric series (using that $N$ is sufficiently large), we obtain
$\operatorname*{{}\mathbb{E}}\!\big[\mathopen{}\mathclose{{\left|\mathbf{z}}}\right|^{2}\big]^{1/2}\leqslant\big(e^{O(d^{2})}H^{-d/2}+e^{O(d)}Hd^{-d}\big)H^{d}.$
Recalling that $H=\lfloor e^{Cd}\rfloor$ we see that, if $C$ is a sufficiently large absolute constant, the right-hand side is $o(H^{d})$ as $N\to\infty$. This establishes ˜6 and concludes the proof of Proposition˜2.2. ∎
Remark 2.5.
For Theorem˜3.18, we obtain (with high probability) an asymptotic formula for quantities similar to ˜4, which is then combined with techniques from geometry of numbers to prove the existence of a short lattice basis.
This requires a more careful analysis of the small-order characters, which may be more numerous than in the special case treated above. Moreover, the analysis of large-order characters is much more delicate, as an exceptional character $\psi$ (in the sense of Lemma˜2.4) may have more representations of the form $\psi=\chi^{h}$ (for some $\chi\in\widehat{G}$ and $|h|<H$) than the previous argument can handle. This issue is resolved by performing Fourier analysis on a well-chosen subgroup of $(\mathbb{Z}/N\mathbb{Z})^{\times}$, introduced in Section˜3.1.
3. Proof of the existence of a short lattice basis
@@ 3.1. Restricting to a convenient subgroup of $\widehat{(\mathbb{Z}/N\mathbb{Z})^{\times}}$
The goal of this preliminary section is to define a subgroup of the character group of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ that does not have too many elements of small order. This will be crucial for Section˜3.3, to reduce the influence of potential counterexamples to the Generalised Riemann Hypothesis. For example, we need to avoid having many characters $\chi_{1},\ldots,\chi_{k}$ modulo $N$ such that $\chi_{i}^{2}=\psi$ for all $i$, where $\psi$ is an exceptional character (in the sense that $L(s,\psi)$ has a zero very close to the line $\operatorname{Re}s=1$). The definition of the suitable subgroup depends on the precise structure of $(\mathbb{Z}/N\mathbb{Z})^{\times}$.
Notation 3.1.
If $G$ is a finite abelian group (written multiplicatively) and $M\geqslant 1$, we write $G^{M}:=\{x^{M}:x\in G\}$ for the subgroup of $M$th powers in GG. [^11] In particular, $G^{M}$ is not the $M$-fold direct product of $G$.
@@
Let $N>2$, $d:=\lceil\!\sqrt{\log N}\rceil$, $G:=(\mathbb{Z}/N\mathbb{Z})^{\times}$ and let ${M_{*}}\geqslant 1$ be the integer defined in the statement of Lemma˜3.4.
@@
which is $\leqslant h^{d/10}$ by Lemma˜3.4. ∎
@@ 3.2. Lattice point counting via characters
To prove Theorem˜3.18, we will need to study the following quantities.
@@
Observe that $\psi\big(b_{i}^{Me_{i}}\big)=\iota(\psi)(b_{i}^{e_{i}})$, where $\iota$ is the isomorphism $\widehat{G^{M}}\xrightarrow{\sim}\widehat{G}^{M}$ defined in Lemma˜3.2. We may thus rewrite this equality as
@@
$\sum_{\begin{subarray}{c}(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\\ \max_{i}|e_{i}|\leqslant H\end{subarray}}\mathbf{1}_{\prod_{i}\mathbf{b}_{i}^{Me_{i}}\equiv 1\,(\mathrm{mod}\,N)}=\frac{1}{\big\lvert\widehat{G}^{M}\big\rvert}\sum_{\chi\in\widehat{G}^{M}}\sum_{\begin{subarray}{c}(e_{1},\ldots,e_{d})\in\mathbb{Z}^{d}\\ \max_{i}|e_{i}|\leqslant H\end{subarray}}\prod_{i=1}^{d}\chi^{e_{i}}(b_{i})=\frac{1}{\big\lvert\widehat{G}^{M}\big\rvert}\sum_{\chi\in\widehat{G}^{M}}F_{\chi,H}(b_{1},\ldots,b_{d})$
@@
It is fairly straightforward to estimate $F_{\chi,H}(b_{1},\ldots,b_{d})$ when $\chi$ is a character of small order in $\widehat{G}$. The contribution of these small-order characters gives the expected main term for ˜9.
@@
where the first equality follows from Lemma˜3.2 and the second from the canonical isomorphism $H_{1}^{\perp}\cong\widehat{G_{1}/H_{1}}$ for any subgroup $H_{1}$ of a finite abelian group $G_{1}$.
Clearly, if $\chi\in\mathopen{}\mathclose{{\left\langle b_{1},\ldots,b_{d}}}\right\rangle^{\perp}\cap\widehat{G}^{M}$, we have $F_{\chi,H}(\underline{b})=(2H+1)^{d}$. Hence, by ˜11, those characters contribute
@@
Note that $C_{I,(z_{i})}$ is either the empty set, or a coset of $\mathopen{}\mathclose{{\left\langle b_{1},\ldots,b_{d}}}\right\rangle^{\perp}\cap\widehat{G}^{M}$ in $\widehat{G}^{M}$. Moreover, if $\chi$ has order at most $e^{10d}$, this set $C_{I,(z_{i})}$ can only be non-empty if all $z_{i}$ are roots of unity of order $\leqslant e^{10d}$, and there are $\leqslant e^{20d}$ such roots of unity. We conclude that
@@
Thus, we can bound the right-hand side of ˜12 by
@@
where we used that $H\geqslant e^{30d}$ in the last step. By ˜11, this means that
@@ 3.3. Bounding the contribution of large-order characters
Proposition˜3.9 is a standard consequence of the explicit formula for $L(s,\chi)$ and is proved in Appendix˜A. The logarithmic factors can be somewhat improved, but such refinements are irrelevant here.
The Generalised Riemann Hypothesis is the claim that, for every Dirichlet character $\chi$, $L(s,\chi)$ has no zero $\rho$ with $\frac{1}{2}<\operatorname{Re}\rho<1$. Assuming GRH, Proposition˜3.9 thus implies almost square-root cancellation for character sums over primes. Proposition˜3.10 will serve as an unconditional substitute for GRH.
@@
This is [jutila, Theorem 1 (1.7)]. ∎
Our goal is to control the total contribution of all large-order characters in ˜10. We will do so in Proposition˜3.14 (restricting to a suitable sugbroup of $\widehat{G}$). We first prove the following $L^{2}$ bound for a single large-order character.
@@
By Proposition˜3.10 (the zero-density estimate), the number of Dirichlet characters modulo $N$ whose associated $L$-function has a zero in the region
(13) $\mathopen{}\mathclose{{\left\{s\in\mathbb{C}\ :\ {1-\frac{1}{10d}}<\operatorname{Re}{s}\leqslant 1,\ \mathopen{}\mathclose{{\left|\operatorname{Im}{s}}}\right|\leqslant{X}}}\right\}$
is bounded by
@@
If a non-principal character $\psi$ modulo $N$ has no zero in the region ˜13, then by Proposition˜3.9 we have
(15) $\mathbb{E}\mathopen{}\mathclose{{\left[\psi(\mathbf{b}_{1})}}\right]\ll X^{-1/(10d)}(\log(NX))^{3}\ll d^{-100}d^{6}\ll d^{-94}.$
@@
We can use ˜15 to bound the term $\mathbb{E}\mathopen{}\mathclose{{\left[\chi^{h}(\mathbf{b}_{1})}}\right]$, unless $\chi^{h}$ is principal or $L(s,\chi^{h})$ has a zero in the region ˜13. By ˜14, the number of values of $h$ with $|h|\leqslant 2H$ for which one of these two situations occurs is
@@
By ˜15, we deduce that, for all but $O(e^{-d}H)$ values of $|h|\leqslant 2H$, the bound $\mathbb{E}\mathopen{}\mathclose{{\left[\chi^{h}(\mathbf{b}_{1})}}\right]\ll d^{-94}$ holds. Hence
$\mathbb{E}\mathopen{}\mathclose{{\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}}\right]\leqslant(2H+1)^{d}\Big(O(e^{-d}H)+O(d^{-94}H)\Big)^{d}\ll e^{O(d)}d^{-94d}H^{2d}\ll d^{-10d}H^{2d},$
@@
Lemma˜3.11 applies to all characters $\chi$ of order at least $e^{10d}$, but gives a relatively weak upper bound. In Lemma˜3.13, we will prove that a stronger bound can be obtained if a small set of exceptions is allowed. We begin by describing the exceptional characters in the following lemma.
@@
Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma˜3.11. Let $H\geqslant e^{10d}$.
@@
Then there are integers $2H\leqslant h_{1},h_{2}\leqslant 2H$ with $0<h_{2}-h_{1}\leqslant e^{3t}$ such that both $L(s,\chi^{h_{1}})$ and $L(s,\chi^{h_{2}})$ have a zero in the region
(16) $\mathopen{}\mathclose{{\left\{s\in\mathbb{C}\ :\ {1-\frac{t}{100d}}<\operatorname{Re}{s}\leqslant 1,\ \mathopen{}\mathclose{{\left|\operatorname{Im}{s}}}\right|\leqslant{X}}}\right\}.$
@@
As in the proof of Lemma˜3.11, we have
@@
Let $I_{1}$ be the set of all $2H\leqslant h\leqslant 2H$ such that $\chi^{h}$ is principal. Since $\chi$ has order $\geqslant e^{10d}$, we have
@@
Let $I_{2}$ be the set of all $2H\leqslant h\leqslant 2H$ such that $L(s,\chi^{h})$ has a zero in the region ˜16. By contradiction, suppose that the conclusion of Lemma˜3.12 does not hold. Then any sub-interval of [2H,2H][-2H,2H] of length $e^{3t}$ contains at most one element of $I_{2}$. This implies that
@@
Moreover, for any integer $h\in[-2H,2H]\setminus(I_{1}\cup I_{2})$, the character $\chi^{h}$ is non-principal and has no zero in the region ˜16, which by Proposition˜3.9 implies that
@@
We wish to use a zero-density estimate again to show that there are few characters satisfying the conclusion of Lemma˜3.12. For this step to work, we need to restrict to the subgroup $\widehat{G}^{{M_{*}}}$ of the full character group $\widehat{G}$, where $M_{*}$} is the integer defined in Lemma˜3.4.
@@
Let $t_{0}\geqslant 1$ be the constant from Lemma˜3.12. Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma˜3.11. Let $H\geqslant e^{10d}$ and let ${M_{*}}\geqslant 1$ be the integer defined in Lemma˜3.4.
@@
(18) $\mathopen{}\mathclose{{\left|\big\{\chi\in\widehat{G}^{{M_{*}}}\ :\ \mathrm{ord}(\chi)\geqslant e^{10d},\ \mathbb{E}\mathopen{}\mathclose{{\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}}\right]>e^{-2td}H^{2d}\big\}}}\right|\ll e^{td/2}.$